a flower bed is in the shape of a triangle with one side twice the length of the shortest side and the third side is 12 ft more than the length of the shortest side the perimeter is 156

Answer by checkley79(3341)

LET X=THE SHORTER SIDE.
LET 2X=THE NEXT SIDE.
LET X+12=THE LONGEST SIDE.
PERIMETER=X+2X+X+12
156=4X+12
156-12=4X
144=4X
X=144/4
36 ANS. FOR THE SHORTEST SIDE.
2*36=72 ANS FOR THE NEXT SIDE.
36+12=48 ANS. FOR THE LONGEST SIDE.
PROOF:
156=36+72+48
156=156